OpenStudy (anonymous):

find power series (maclaurin series): integral [ ln(1+x^2) ] dx

OpenStudy (anonymous):

What are the limits on the integral? if any

OpenStudy (anonymous):

Ah just realized you don't need to have a definite integral... Recall the power series for \(\ln(1+x)\): \[\ln(1+x)=\sum_{k=1}^\infty \frac{(-1)^{k+1}x^k}{k}\quad(|x|<1)\] Replace \(x\) with \(x^2\), then take the antiderivative of the series.

OpenStudy (anonymous):

yeah... i was hoping someone would actually do the dirty work though and show me

OpenStudy (anonymous):

You and most of everyone else